OpenStudy (anonymous):

Find the roots of the function f(x)=(x-8)^2-1

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

i have no clue honestly

OpenStudy (triciaal):

to find the root set the function = 0 and solve for x

OpenStudy (anonymous):

\[(x-8)^{2}-1 = x^2-16x+64-1=x^{2}-16x+63 = 0\]

OpenStudy (anonymous):

ok so i would do (x-8)^2-1=0?

OpenStudy (anonymous):

Follow ginochen

OpenStudy (anonymous):

Solve for \( \large x^{2}-16x+63 = 0\)

OpenStudy (triciaal):